Confidence Level 99%
SUMMARY STATISTICS
n 65
Mean 16.000
Std Dev 1.400
RESULTS
SE
Z
Margin of Error
Lower Limit
Upper Limit
Answer #1

Given CI level is 99%, hence α = 1 - 0.99 = 0.01
α/2 = 0.01/2 = 0.005, Zc = Z(α/2) = 2.58


SE = σ/sqrt(n)
= 1.4/sqrt(65)
= 0.1736

z value at 99% = 2.58

ME = zc * σ/sqrt(n)
ME = 2.58 * 1.4/sqrt(65)
ME = 0.4480

CI = (xbar - Zc * s/sqrt(n) , xbar + Zc * s/sqrt(n))
CI = (16 - 2.58 * 1.4/sqrt(65) , 16 + 2.58 * 1.4/sqrt(65))
CI = (15.552 , 16.448)

Lower limit = 15.5520
Upper limit = 16.4480